APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday January 15, 2002 (MJD 52290) ---=== OBSERVING TEAM ===--- Mike Harvanek (APO) Jurek Krzesinski (APO) Pete Newman (APO; swing) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Heavy clouds all night. Never opened. ---=== OBSERVING LOG ===--- It was totally overcast with heavy clouds at sunset. Since we could not take data in these conditions we decided to wait for the clouds to clear before removing the enclosure. While waiting, we saw the following error in watcher: altitude (Windscreen touched in altitude: UP (Jan 15 19:10:51): Jan 15 19:11:03 MST ) When this occurred, the telescope had been motionless for about 2 hours. The telescope was in stow position with 2 stop buttons in, the spectro auto fill was connected, and all amplifiers and the wind baffle were turned on. The east door was fully opened and the west door was cracked open about a foot. There were no high speed wind gusts that we know of. We're not sure if this was a false alarm or if something slipped but either way, we thought it was very unusual. The clouds never cleared and we never opened. ---=== IMAGING RUN SUMMARY ===--- Run Stripe Flavor Lambda Last Begin End Frame Comments ------------------------------------------------------------------ 2890 100 O ignore 166.79 171.00 42 ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot --------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields nGood rowMean rowSig colMean colSig rot --------------------------------------------------------------- ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 12688, plate -9999 ------- 22:54 12688 0.0 bias {{}} 23:01 12689 0.0 bias {{}} 23:23 12690 0.0 bias {{}} ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument getclock Az Alt Rot Scale ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: none Stays: none ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 16:44 37 4 10 216 (SW) 454 - mcl 17:14 37 5 13 225 (SW) 347 - mcl 17:44 36 5 16 220 (SW) 497 - tcl 18:16 36 6 9 255 (WSW) 494 - tcl 18:49 37 3 12 257 (WSW) 297 - tcl 19:21 38 2 10 264 (W) 196 - tcl 19:53 38 2 15 255 (WSW) 189 - tcl 20:25 37 3 20 231 (SW) 186 - tcl 20:57 36 4 16 237 (WSW) 229 - tcl 21:30 36 6 18 232 (SW) 332 - tcl 22:02 36 6 18 260 (W) 282 - tcl 22:34 36 7 16 243 (WSW) 277 - tcl 23:07 36 6 16 252 (WSW) 300 - tcl 23:39 35 6 19 256 (WSW) 340 - tcl 00:11 35 7 13 265 (W) 341 - tcl 00:43 34 7 17 247 (WSW) 315 - tcl 01:16 34 8 12 246 (WSW) 338 - tcl 01:48 33 8 18 245 (WSW) 285 - tcl 02:18 33 7 16 259 (W) 221 - tcl 02:48 33 7 18 245 (WSW) 189 - tcl 03:19 33 6 17 253 (WSW) 197 - tcl 03:51 32 8 16 250 (WSW) 175 - tcl 04:24 32 8 13 251 (WSW) 216 - tcl 04:57 32 9 16 238 (WSW) 184 - tcl [pmt]cl = partly, mostly, totally cloudy ---=== TELESCOPE STATUS ===--- 23:30 Doors open, fans on. 12:00 Doors closed, fans off. Never opened. Telescope stowed at 30 degrees with cartridge 4 (plate 807) mounted. Counterweights at 270. Spectro auto-fill connected. Imager dewar 12:11 188 lb Spectro dewar 12:11 90 lb Building interlocks bypassed. Alt, az slip detection interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_64_3c Watcher: v2_13_2a MCP: v5_8_7 TPM: v2_4_5 AstroDa: v14_40 TCC: v2.5.5 sdssProcedures: v1_37 TccMon: v1_8 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0740 0.0750 Axial B 0.8070 0.8020 Axial C 0.0750 0.0710 Trans D -1.5000 -1.4860 Lateral E 1.8220 1.8540 Lateral F 1.4070 1.4240 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 2.8 Alt.: 30.369169 MIGS TONIGHT NOMINAL Axial A 1.5440 1.5500 Axial B 1.2200 1.2260 Axial C 1.2440 1.2480 Trans D 0.2040 0.2320 GALILS Commanded: 1599636. 1565299. 1586389. -10395. -9766. Actual: 1599620. 1565254. 1586350. -10395. -9766. SETMIR VALUES SecDesOrient: 1257.00 -3.40 -10.00 0.00 133.82 SecOrient: 1257.03 -3.40 -9.99 0.01 133.82 ---=== PROBLEMS IN DETAIL ===--- Windscreen touched telescope when both were supposedly motionless: ------------------------------------------------------------------ altitude (Windscreen touched in altitude: UP (Jan 15 19:10:51): Jan 15 19:11:03 MST ) When this occurred, the telescope had been motionless for about 2 hours. The telescope was in stow position with 2 stop buttons in, the spectro auto fill was connected, and all amplifiers and the wind baffle were turned on. The east door was fully opened and the west door was cracked open about a foot. There were no high speed wind gusts that we know of. We're not sure if this was a false alarm or if something slipped but either way, we thought it was very unusual.